A mfg. concern, which has adopted standard costing, furnishes the following information:

Standard:

Material for 70 kg. Of finished products ----> 100 kgs.

Price of materials . -----> Rs. 1 per kg

Actual:

Output ----> 210,000 kgs

Material used -----> 280,000 kgs.

Cost of materials  ----> Rs. 2,52,000

Calculate material cost variance, material price variance, and material usage variance?


1
Expert's answer
2022-01-14T10:18:02-0500

Material cost Variance= "(Standard Quantity\times Standard Price)- (Actual Quantity\times Actual Price)"

= "(Standard Quantity\times Standard Price)- (Actual Quantity\times Actual Price)"

Standard quantity="\frac {(2100000\times100)} {70}=3,000,000"

Actual price= "\frac {252000} {280000}=0.9"

MCV="(3000000\times 1)- (280000\times 0.9)"

"(3000000)- (252000)= 2,748,000 (F)"

Material Price Variance="(Budgeted Price-Actual Price)\times Quantity of material Used"


= "(1-0.9)\times 280,0000"

= 28,000 (F)


Material Usage Variance

"MUV=(Standard Quantity allowed-Actual quantity used)\times standard price per unit of material"

"=(3000000-280,000)\times 1"

"= 2,720,000"